Grade 12Mechanics

How is the time period of a spring mass system = 2 pi (disp/accn)1/2 ?

Chirag Jain
10 Years agoGrade 12

1 Answer

arun
ApprovedApproved Tutor Answer10 Years ago
standard equation of S.H.M is 
\frac{\partial^2 x}{\partial t^2} + \omega ^2 x =0
a = -\omega ^2 x
\left | a\div x \right |= \omega ^2
2\pi \div T=\omega
T = 2\pi \div \omega
T = 2\pi \div \sqrt{\omega ^2}
T = 2\pi \div \sqrt{a\div x}
T = 2\pi \sqrt{x\div a}
this means time period of of S.H.M is square root of displacement divided by acceleration.